what are the coordinates of the focus and the equation of the directrix for the parabola with equation x=8y^2?
Answer by lwsshak3(11628)

what are the coordinates of the focus and the equation of the directrix for the parabola with equation x=8y^2?
***
y^2=(1/8)x
This is an equation of a parabola that opens right with vertex at the origin(0,0)
Its basic form of equation: y^2=4px.
For given parabola:
axis of symmetry: y=0
4p=1/8
p=1/32
focus: (1/8,0)
directrix: x=-1/8